# A Survey on Multi-Generative Agent System: Recent Advances and New Frontiers

### v1

@@ Abstract
MGAS have become a research hotspot since the rise of large language models (LLMs). However, with the continuous influx of new related works, the existing reviews struggle to capture them comprehensively. This paper presents a comprehensive survey of these studies. We first discuss the definition of multi-generative agent system (MGAS), a framework encompassing much of previous work. We provide an overview of the various applications of MGAS in (i) solving complex tasks, (ii) simulating specific scenarios, and (iii) evaluating generative agents . Building on previous studies, we also highlight several challenges and propose future directions for research in this field.
A Survey on Multi-Generative Agent System:
Recent Advances and New Frontiers
Shuaihang Chen  Yuanxing Liu  Wei Han  Weinan Zhang [^2] Corresponding author.  Ting Liu Research Center for Social Computing and Information Retrieval Harbin Institute of Technology, China {shchen, yxliu, whan, wnzhang, tliu}@ir.hit.edu.cn
@@ 1 Introduction
MAS has seen significant expansion owing to its adaptability and ability to address complex, distributed challenges Balaji and Srinivasan (2010). Compared to single-agent settings Gronauer and Diepold (2022), multi-agent system (MAS) provides a more accurate representation of the real world, as many real-world applications naturally involve multiple decision-makers interacting simultaneously. However, constrained by traditional  reinforcement learning (RL) agent parameters and the absence of general knowledge and capabilities, agents are unable to tackle complex decision-making tasks, such as collaborating with other agents for the development Qian et al. (2024b). In recent years, large language models (LLMs), e.g. Llama 3 Dubey et al. (2024), and GPT-4 OpenAI et al. (2024), have achieved notable successes, training on massive web corpus Radford et al. . Compared with RL, generative agents, with large language model (LLM) as the core control agents, can be better at reasoning, long-trajectory decision-making, etc., even without training Shinn et al. (2023). Furthermore, generative agents offer natural language interfaces for interacting with humans, making these interactions more flexible and easier to explain Park et al. (2023).
@@
Based on these advantages, multi-generative agent system (MGAS) emerged. Researchers have surveyed these emerging works and proposed a general framework Guo et al. (2024). However, as the number of related studies continues to grow, some works have emerged that fall outside the scope of the original framework. In this paper, we provide a new perspective based on previous reviews of multi-generative agent systems (MGASs) with a focus on recent advancements and discuss potential research directions. We collected 125 papers published in top artificial intelligence conferences, such as *ACL, NeurIPS, AAAI, and ICLR, in 2023 and 2024, along with some unpublished yet valuable papers from arXiv. [^1] The list of papers included in this survey can be found in https://github.com/bianhua-12/Multi-generative_Agent_System_survey Based on the purpose of application on MGAS, we summarize the application of MGAS as task-solving, simulation for specific problems, and evaluation of generative agents. Figure 1 illustrates the framework we propose for MGAS application. (i) Solving complex tasks. Multi-agents will naturally split tasks into subtasks, which will improve task performance. (ii) Simulating for specific scenarios. Researchers see MGAS as a sandbox for simulating problems in a specific domain. (iii) Evaluating generative agents. Compared with traditional task evaluation, MGAS has the capability of dynamic assessment, which is more flexible and harder for data leakage. For each category, we will discuss representative MGAS, resources and their evaluation.
Figure 1: Application framework of MGAS and relationship of MGAS, generative agent, and LLM.
@@
Compared to the previous survey Wang et al. (2024a); Zhao et al. (2024c); Chuang and Rogers (2023); Guo et al. (2024); Gao et al. (2023a); Gronauer and Diepold (2022), this survey has the following distinctive contributions: (i) A Taxonomy with greater relevance to current trends : we introduce a more recent taxonomy (shown in Figure 1) based on the purpose of the application of MGAS. (ii) More Resources: we analyze open-source frameworks and research works with benchmarks or datasets to facilitate the research community. (iii) Challege and Future: we discuss the challenges in MGAS, and shed light on future research.
2 Core Components of MGAS
@@
MGAS refers to a system that includes a collection of generative agents capable of interacting and collaborating within a shared environmental setting Wang et al. (2024c). We will discuss generative agents and the environment in the following.
@@ 2.1 Generative Agents
Generative agents refer to the components of MGAS that have role definitions, can perceive the environment, make decisions, and perform complex actions to change the environment Wang et al. (2024a). They can be a player in a game or a user on social media and have the role of driving the development of MGAS and influencing its results.
@@ 2.2 Environment
3 MGASs for Solving Complex Tasks
In most scenarios in reality, completing a task requires multiple roles, multiple steps, and so on. This is difficult for a single agent, but multiple agents working together will be well suited to this task Islam et al. (2024). Further, each of these agents can be trained independently Shen et al. (2024); Yu et al. (2024). Compared with a single agent, MGAS can achieve better results. That is, the multi-agent collaboration will improve the overall performance Du et al. (2023).
3.1 Representative MGASs for Solving Complex Tasks
@@
MGAS reasoning framework. Multi-stage cooperation is early work. Du et al. (2023) propose a multi-agent debate framework to improve factual correctness and reasoning accuracy, which exceeds the level of a single model. As a representative framework in the open source community, ChatDev Qian et al. (2024b) is a multi-role framework to generate code. Scaling law in agent cooperation is also explored Qian et al. (2024c), finding that there is no significant pattern. Collective Decision-Making refers to agents working independently and voting for the result. Wang et al. (2024d) propose a framework that synchronizes the various agents, and Zhao et al. (2024c) propose an electoral framework to improve reasoning capabilities. Self-Refine refers to the mechanism of self-reflection in MGAS. Researchers propose a framework for applying multi-agents to the natural sciences Chen et al. (2024a) to enhance data analysis, model simulations, and decision-making processes Yin et al. (2024). Zhang et al. (2023a) propose a framework to achieve self-adaptation and adaptive cooperation.
MGAS communication optimization. The fully connected communication in MGAS can lead to issues such as combinatorial explosion and privacy disclosure. Liu et al. (2024a) propose a distributed framework for generative agents to communicate and solve tasks. Zhang et al. (2023a) enhance the cooperation of agents in tasks. DroidSpeak Liu et al. (2024b) uses non-verbal communication, like E-cache or KV-cache, to speed up multi-agent interaction.
3.2 Resources of MGAS for Solving Complex Tasks
We summarize common and open-source MGAS for simulation in Table 1, including code, dataset, and benchmark.
Table 1: Codes and Benchmarks in MGAS for solving tasks studies. “No Code” or “No Benchmark” means the code or benchmark is unavailable.
@@ 3.2 Resources of MGAS for Solving Complex Tasks
Open source community. The open-source and industrial communities have also contributed significantly to the development of MGAS. MetaGPT Hong et al. (2023) assigns different roles to generative agents to form a collaborative entity for complex tasks. Gao et al. (2024) propose AgentScope with message exchange as its core communication mechanism. In the meantime, this work develops a distribution framework that facilitates seamless switching between local and distributed deployments and automatic parallel optimization with minimal effort. Open AI proposes Swarm Ope (2024), an experimental multi-agent orchestration framework that is ergonomic and lightweight. Unlike the previously released Assistants API, Swarm gives developers fine-grained control over context, steps, and tool calls rather than being hosted.
3.3 Evaluation of MGAS for solving complex task
Performance on specific tasks. Shown as Table 1, the performance of MGAS can be evaluated by specific tasks, which is intuitive and convenient. For example, in an APP system Zhang et al. (2023b), the average number of steps and tools used by an agent to complete a specific task are considered as indicators; in BOLAA Liu et al. (2023c), the recall and QA accuracy of intelligent physical examination retrieval are also considered as evaluation indicators; in the Werewolf game Xu et al. (2023c), the win rate of virtual players is naturally also an evaluation indicator; in the job fair system Li et al. (2023) , the proportion of correctly recruited target job seekers by the recruiting party is also an evaluation indicator; in the auction system Chen et al. (2024c), the Spearman correlation coefficient between the predicted and actual prices of goods, as well as the skills of bidders, are also measured by TrueSkill scores Graepel et al. (2007); in Stanford Town Park et al. (2023), the quality of behaviors generated by virtual agents and human agents is manually sorted and evaluated using TrueSkill; in urban simulation systems Xu et al. (2023a), the success rate of completing specific tasks such as navigation is also an evaluation metric.
@@ 3.3 Evaluation of MGAS for solving complex task
4 MGASs for Simulating Specific Scenarios
This section will illustrate the application for MGAS in simulation. Researchers apply agents to simulate a certain scenario to study its impact on a specific subject like social science. On the one hand, compared with rule-based methods Chuang and Rogers (2023), generative agents with natural language communication can be more intuitive for humans. On the other hand, environment determines the properties of the simulation, which is the core of the entire simulation.
4.1 Representative MGASs for Simulating Specific Scenarios
The typical scenarios for MGASs simulations are described as follows. We will introduce the following work according to the subject.
Social domain. Social large-scale experiments in the real world have high costs, and the sheer scale of social participation can sometimes escalate into violence and destruction, posing potential ramifications Mou et al. (2024). Therefore, it is necessary to simulate in the virtual environment; simulation can solve the problem of excessive overhead in the real environment and can simulate the process in the real world for a long time at a faster speed Li et al. (2024a). At the same time, the whole process can be easily repeated, which is conducive to further research. Researchers have done a lot of work to simulate social media scenarios. Based on the social media simulation archetype Park et al. (2022), Park et al. (2023) propose Stanford Town, which leads to a one-day simulation of the life of 25 agents with different occupations in a small American town. At the same time, there was work on emotional propagation influence Gao et al. (2023b), information cocoon room based on recommendation scenario research Wang et al. (2024b), and study of social movements Mou et al. (2024). Researchers propose Urban Generative Intelligence (UGI) Xu et al. (2023a) to address specific urban issues and simulate complex urban systems, providing a multidisciplinary approach to understanding and managing urban complexity. Li et al. (2024a) study doctor agent evolution method by hospital simulation. Because doctor agent training is both inexpensive and highly effective, this work can quickly scale up the agent to handle tens of thousands of cases in just a few days, a task that would take a human doctor years to complete. Pan et al. (2024) propose a huge scale of agent simulation, increasing the number of agents to $10^{6}$. In social game,like Werewolf Xu et al. (2024) , Avalon Lan et al. (2024) , and Minecraft Gong et al. (2024) for MGAS simulation are attempted. Further, some game companies like Netease are also actively experimenting with MGAS in their games.
@@ 4.1 Representative MGASs for Simulating Specific Scenarios
4.2 Resources for MGAS simulation
We summarize common and open-source MGAS for simulation in Table 2, including code and benchmark.
To prove the effectiveness of the simulation, that is, to fit the reality, researchers usually evaluate the simulation system by simulating real data. Therefore, a realistic dataset with dense users and records is very important for evaluation simulation Mou et al. (2024). An ideal dataset will be dense: that is, data with a smaller number of users on the same scale can better evaluate the simulation capability of the MGAS.
@@ 4.2 Resources for MGAS simulation
Table 2: Codes and Benchmarks in MGAS for simulation studies. “No Code” or “No Benchmark” means the code or benchmark is unavailable.
@@
4.3 Evaluation of MGAS simulation
We will discuss the evaluation based on indicators used for assessing MGAS as a whole, rather than the capabilities of individual agents.
Consistency. MGASs necessitate a robust congruence with the real world to ensure the derivation of meaningful and insightful experimental outcomes. In the context of simulation systems, exemplified by UGI Xu et al. (2023a), the primary objective lies in faithfully replicating specific real-world scenarios. When employed for training agents like SMART Yue et al. (2024), only those agents that have undergone rigorous training within a virtual environment that closely mirrors the real environment can be deemed suitable for deployment in real-world settings. Similarly, when utilized for evaluation purposes, such as in AgentSims Lin et al. (2023), the attainment of authentic and reliable evaluation results is contingent upon the virtual environment maintaining a high degree of consistency with its real-world counterpart. Finally, in the system for collecting data such as BOLAA Liu et al. (2023c), consistency also ensures the validity of the data. Therefore, an important performance measure of MGAS is its consistency with the real situation.
@@ 4.3 Evaluation of MGAS simulation
Other specific attributes. It may be difficult to compare the distribution of system behavior directly, but the distribution of some system attributes can be easier to analyze statistically. In a virtual social media platform Abdelzaher et al. (2020), the categories, quantities, average start times, and durations of events occurring in virtual systems and the real world were compared; the auction system Chen et al. (2024c) studied the kernel density estimation map of commodity transaction prices; the virtual environment of the recommendation system Wang et al. (2024b) compared the distribution of user ratings and favorite categories of products, as well as the activity level of agents; in the context of the Werewolf Killing mission, researchers Xu et al. (2023c) compared the behavior distribution of agent players in specific scenarios.
5 MGASs for Evaluating Generative Agents
With LLMs prevailing in the community, how to evaluate the ability of LLMs is an open question. Existing evaluation methods suffer from the following shortcomings: (i) constrained evaluation abilities, (ii) vulnerable benchmarks, and (iii) unobjective metrics. The complexity and diversity of MGAS have indicated that MGAS can evaluate LLMs. However, how to design specific evaluation indicators and evaluation methods has puzzled researchers. Similarly, MGAS can also be used in training generative agents. We summarize three aspects of training: (i) Supervised Fine-Tuning (SFT) (ii) reinforcement learning (RL) (iii) Synthesizing data for training.
5.1 Representative MGASs for Evaluating Generative Agents
MGAS can provide rewards to agents, and these rewards can be used to evaluate or train generative agents, which will be discussed below.
Evaluation of generative agents. Researchers study generative agents by putting them into MGAS. In MGAS, researchers can further study the LLM’s strategic capabilities in different scenes, such as long strategic ability Chen et al. (2024c), leadership strategy Xu et al. (2023c) and competitiveness strategy Zhao et al. (2024b). In the emotional field, MuMA-ToM Shi et al. (2024) is used to evaluate the ability of agents to understand and reason about human interactions in a real home environment through video and text descriptions.
Training on generative agents. Li et al. (2024c) enhance the data to Supervised Fine-Tuning (SFT) generative agents with MGASXu et al. (2023c) have created generative agents to overcome the intrinsic bias from LLMs by proposing a novel framework that powers generative agents with multi-agent reinforcement learning Xu et al. (2023c). For MGAS, Yue et al. (2024) split complex trajectories in knowledge-intensive tasks into subtasks, proposing a co-training paradigm of the multi-agent framework, Long- and Short-Trajectory Learning, which ensures synergy while keeping the fine-grained performance of each agent. RLHF has been criticized for its high cost. Liu et al. (2023a) propose an alignment scheme based on a multi-agent system, effectively addressing instability and reward gaming concerns associated with reward-based RL optimization. Either way, the MGAS is essentially viewed as an environment in RL with different ways of getting rewards from the environment.
5.2 Resources of MGAS for evaluations

### v2

# A Survey on LLM-based Multi-Agent System: Recent Advances and New Frontiers in Application

@@ Abstract
LLM-MAS have become a research hotspot since the rise of large language models (LLMs). However, with the continuous influx of new related works, the existing reviews struggle to capture them comprehensively. This paper presents a comprehensive survey of these studies. We first discuss the definition of LLM-based Multi-Agent Systems (LLM-MAS), a framework encompassing much of previous work. We provide an overview of the various applications of LLM-MAS in (i) solving complex tasks, (ii) simulating specific scenarios, and (iii) evaluating generative agents. Building on previous studies, we also highlight several challenges and propose future directions for research in this field.
A Survey on LLM-based Multi-Agent System:
Recent Advances and New Frontiers in Application
Shuaihang Chen1  Yuanxing Liu1  Wei Han1  Weinan Zhang1222Corresponding author.  Ting Liu1 1Research Center for Social Computing and Information Retrieval Harbin Institute of Technology, China {shchen, yxliu, whan, wnzhang, tliu}@ir.hit.edu.cn
@@ 1 Introduction
MAS have seen significant expansion owing to its adaptability and ability to address complex, distributed challenges Balaji and Srinivasan (2010). Compared to single-agent settings Gronauer and Diepold (2022), Multi-Agent Systems (MAS) provide a more accurate representation of the real world, as many real-world applications naturally involve multiple decision-makers interacting simultaneously. However, constrained by traditional  reinforcement learning (RL) agent parameters and the absence of general knowledge and capabilities, agents are unable to tackle complex decision-making tasks, such as collaborating with other agents for the development Qian et al. (2024b). In recent years, large language models (LLMs), e.g. Llama 3 Dubey et al. (2024), and GPT-4 OpenAI et al. (2024), have achieved notable successes, training on a massive web corpus Radford et al. . Compared with RL, generative agents, with large language model (LLM) as the core control agents, can be better at reasoning, long-trajectory decision-making, etc., even without training Shinn et al. (2023). Furthermore, generative agents offer natural language interfaces for interacting with humans, making these interactions more flexible and easier to explain Park et al. (2023). Based on these advantages, LLM-based Multi-Agent Systems (LLM-MAS) emerged. Researchers have surveyed these emerging works and proposed a general framework Guo et al. (2024). However, as the number of related studies continues to grow, some works have emerged that fall outside the scope of the original framework. In this paper, we provide a new perspective based on previous reviews of LLM-based Multi-Agent Systems (LLM-MAS) with a focus on recent advancements and discuss potential research directions. We collected 125 papers published in top artificial intelligence conferences, such as *ACL, NeurIPS, AAAI, and ICLR, in 2023 and 2024, along with some unpublished yet valuable papers from arXiv. [^1] The list of papers included in this survey can be found in https://github.com/bianhua-12/Multi-generative_Agent_System_survey Based on the purpose of LLM-MAS, we summarize the application of LLM-MAS as task-solving, simulation for specific problems, and evaluation of generative agents. Figure 1 illustrates the framework we propose for LLM-MAS application. (i) Solving complex tasks. Multi-agents will naturally split tasks into subtasks, which will improve task performance. (ii) Simulating for specific scenarios. Researchers see LLM-MAS as a sandbox for simulating problems in a specific domain. (iii) Evaluating generative agents. Compared with traditional task evaluation, LLM-MAS has the capability of dynamic assessment, which is more flexible and harder for data leakage. For each category, we will discuss representative LLM-MAS, resources, and their evaluation.
Figure 1: Overview of the application framework and relationship of LLM-MAS, generative agent, and LLM. Dashed-bordered right-angled rectangles represent content aligned with previous surveys, while rounded rectangles indicate original contributions introduced in this study.
@@
Compared to the previous survey Guo et al. (2024); Li et al. (2024d); Han et al. (2024); Gronauer and Diepold (2022), this survey has the following distinctive contributions: (i) A Taxonomy focusing on application of LLM-MAS: we introduce a more recent taxonomy (taxonomy and difference are shown in Figure 1) based on the purpose of the application of LLM-MAS. (ii) More Resources: we analyze open-source frameworks and research works with benchmarks or datasets to facilitate the research community. (iii) Challenges and Future: we discuss the challenges in LLM-MAS, and shed light on future research.
2 Core Components of LLM-MAS
@@
LLM-MAS refer to a system that includes a collection of generative agents capable of interacting and collaborating within a shared environmental setting Wang et al. (2024c). We will discuss generative agents and the environment in the following.
@@ 2.1 Generative Agents
Generative agents refer to the components of LLM-MAS that have role definitions, can perceive the environment, make decisions, and perform complex actions to change the environment Wang et al. (2024a). They can be a player in a game or a user on social media and have the role of driving the development of LLM-MAS and influencing its results.
@@ 2.2 Environment
3 LLM-MAS for Solving Complex Tasks
Completing a complex task usually requires multiple roles, multiple steps, and so on. This is difficult for a single agent, but multiple agents working together will be well suited to this task Islam et al. (2024). Further, each of these agents can be trained independently Shen et al. (2024); Yu et al. (2024). Compared with a single agent, LLM-MAS can achieve better results. That is, the multi-agent collaboration will improve the overall performance Du et al. (2023).
3.1 Representative LLM-MAS for Solving Complex Tasks
@@
LLM-MAS reasoning framework. We summarize three aspects by the pipeline of reasoning, including: (i) multi-stage framework, (ii) collective decision-making framework, and (iii) self-refine framework. That is, the multi-stage framework refers to a pipeline where agents act as serial problem solvers at different stages Qian et al. (2024b), while collective decision-making Zhao et al. (2024c) refers to different agents voting or debating for one goal. Self-Refine refers to the mechanism of self-reflection in LLM-MAS. Researchers propose a framework for applying multi-agents to the natural sciences Chen et al. (2024a) to enhance data analysis, model simulations, and decision-making processes Yin et al. (2024). Zhang et al. (2023a) propose a framework to achieve self-adaptation and adaptive cooperation. Scaling law in agent cooperation is also explored Qian et al. (2024c), finding that there is no significant pattern.
LLM-MAS communication optimization. The fully connected communication in LLM-MAS can lead to issues such as combinatorial explosion and privacy disclosure. Based on this, we summarize two aspects in Communication Optimization, including: (i) speed optimization and (ii) distributed discussion. Speed optimization refers to researchers trying to speed up the communication of agents, for example, with non-verbal communication Liu et al. (2024b) or shorter generation Chen et al. (2024g). While distributed discussion refers to agents trying to solve tasks without enough information Liu et al. (2024a). Agents need to communicate with each other to achieve their goals Zhang et al. (2023a), even without complete information in one agentLiu et al. (2024a).
3.2 Resources of LLM-MAS for Solving Complex Tasks
We summarize common and open-source LLM-MAS for simulation in Table 1, including code, dataset, and benchmark.
Table 1: Codes and Benchmarks in LLM-MAS for solving tasks studies. “No Code” or “No Benchmark” means the code or benchmark is unavailable.
@@ 3.2 Resources of LLM-MAS for Solving Complex Tasks
Open source community. The open-source and industrial communities have also contributed significantly to the development of LLM-MAS. MetaGPT Hong et al. (2023) assigns different roles to generative agents to form a collaborative entity for complex tasks. Gao et al. (2024) propose AgentScope with message exchange as its core communication mechanism. In the meantime, this work develops a distribution framework that facilitates seamless switching between local and distributed deployments and automatic parallel optimization with minimal effort. Open AI proposes Swarm Ope (2024), an experimental multi-agent orchestration framework that is ergonomic and lightweight. Unlike the previously released Assistants API, Swarm gives developers fine-grained control over context, steps, and tool calls rather than being hosted.
3.3 Evaluation of LLM-MAS for solving complex task
Performance on specific tasks. Shown as Table 1, the performance of LLM-MAS can be evaluated by specific tasks, which is intuitive and convenient. For example, in an APP system Zhang et al. (2023b), the average number of steps and tools used by an agent to complete a specific task are considered as indicators; in BOLAA Liu et al. (2023c), the recall and QA accuracy of intelligent physical examination retrieval are also considered as evaluation indicators; in the Werewolf game Xu et al. (2023c), the win rate of virtual players is naturally also an evaluation indicator; in the job fair system Li et al. (2023) , the proportion of correctly recruited target job seekers by the recruiting party is also an evaluation indicator; in the auction system Chen et al. (2024c), the Spearman correlation coefficient between the predicted and actual prices of goods, as well as the skills of bidders, are also measured by TrueSkill scores Graepel et al. (2007); in Stanford Town Park et al. (2023), the quality of behaviors generated by virtual agents and human agents is manually sorted and evaluated using TrueSkill; in urban simulation systems Xu et al. (2023a), the success rate of completing specific tasks such as navigation is also an evaluation metric.
@@ 3.3 Evaluation of LLM-MAS for solving complex task
4 LLM-MAS for Simulating Specific Scenarios
This section will illustrate the application for LLM-MAS in simulation. Researchers apply agents to simulate a certain scenario to study its impact on a specific subject like social science. On the one hand, compared with rule-based methods Chuang and Rogers (2023), generative agents with natural language communication can be more intuitive for humans. On the other hand, environment determines the properties of the simulation, which is the core of the entire simulation.
4.1 Representative LLM-MAS for Simulating Specific Scenarios
The typical scenarios for LLM-MAS simulations are described as follows. We will introduce the following work according to the subject.
Social domain. Social large-scale experiments in the real world have high costs, and the sheer scale of social participation can sometimes escalate into violence and destruction, posing potential ramifications Mou et al. (2024). Therefore, it is necessary to simulate in the virtual environment; simulation can solve the problem of excessive overhead in the real environment and can simulate the process in the real world for a long time at a faster speed Li et al. (2024a). At the same time, the whole process can be easily repeated, which is conducive to further research. Researchers have done a lot of work to simulate social media scenarios. Based on the social media simulation archetype Park et al. (2022), Park et al. (2023) propose Stanford Town, which leads to a one-day simulation of the life of 25 agents with different occupations in a small American town. At the same time, there was work on emotional propagation influence Gao et al. (2023b), information cocoon room based on recommendation scenario research Wang et al. (2024b), and study of social movements Mou et al. (2024). Researchers propose Urban Generative Intelligence (UGI) Xu et al. (2023a) to address specific urban issues and simulate complex urban systems, providing a multidisciplinary approach to understanding and managing urban complexity. Li et al. (2024a) study doctor agent evolution method by hospital simulation. Because doctor agent training is both inexpensive and highly effective, this work can quickly scale up the agent to handle tens of thousands of cases in just a few days, a task that would take a human doctor years to complete. Pan et al. (2024) propose a huge scale of agent simulation, increasing the number of agents to $10^{6}$. In social game,like Werewolf Xu et al. (2024) , Avalon Lan et al. (2024) , and Minecraft Gong et al. (2024) for LLM-MAS simulation are attempted. Further, some game companies like Netease are also actively experimenting with LLM-MAS in their games.
@@ 4.1 Representative LLM-MAS for Simulating Specific Scenarios
4.2 Resources for LLM-MAS simulation
We summarize common and open-source LLM-MAS for simulation in Table 2, including code and benchmarks.
To prove the effectiveness of the simulation, that is, to fit the reality, researchers usually evaluate the simulation system by simulating real data. Therefore, a realistic dataset with dense users and records is very important for evaluation simulation Mou et al. (2024). An ideal dataset will be dense: that is, data with a smaller number of users on the same scale can better evaluate the simulation capability of the LLM-MAS.
@@ 4.2 Resources for LLM-MAS simulation
Table 2: Codes and Benchmarks in LLM-MAS for simulation studies. “No Code” or “No Benchmark” means the code or benchmark is unavailable.
@@
4.3 Evaluation of LLM-MAS simulation
We will discuss the evaluation based on indicators used for assessing LLM-MAS as a whole, rather than the capabilities of individual agents.
Consistency. LLM-MAS necessitate a robust congruence with the real world to ensure the derivation of meaningful and insightful experimental outcomes. In the context of simulation systems, exemplified by UGI Xu et al. (2023a), the primary objective lies in faithfully replicating specific real-world scenarios. When employed for training agents like SMART Yue et al. (2024), only those agents that have undergone rigorous training within a virtual environment that closely mirrors the real environment can be deemed suitable for deployment in real-world settings. Similarly, when utilized for evaluation purposes, such as in AgentSims Lin et al. (2023), the attainment of authentic and reliable evaluation results is contingent upon the virtual environment maintaining a high degree of consistency with its real-world counterpart. Finally, in the system for collecting data such as BOLAA Liu et al. (2023c), consistency also ensures the validity of the data. Therefore, an important performance measure of LLM-MAS is its consistency with the real situation.
@@ 4.3 Evaluation of LLM-MAS simulation
5 LLM-MAS for Evaluating Generative Agents
With LLMs prevailing in the community, how to evaluate the ability of LLMs is an open question. Existing evaluation methods suffer from the following shortcomings: (i) constrained evaluation abilities, (ii) vulnerable benchmarks, and (iii) unobjective metrics. The complexity and diversity of LLM-MAS have indicated that LLM-MAS can evaluate LLMs. However, how to design specific evaluation indicators and evaluation methods has puzzled researchers. Similarly, LLM-MAS can also be used in training generative agents. We summarize three aspects of training: (i) Supervised Fine-Tuning (SFT) (ii) reinforcement learning (RL) (iii) Synthesizing data for training.
5.1 Representative LLM-MAS for Evaluating Generative Agents
LLM-MAS can provide rewards to agents, and these rewards can be used to evaluate or train generative agents, which will be discussed below.
Evaluation of generative agents. Researchers study generative agents by putting them into LLM-MAS. In LLM-MAS, researchers can further study the LLM’s strategic capabilities in different scenes, such as long strategic ability Chen et al. (2024c), leadership strategy Xu et al. (2023c) and competitiveness strategy Zhao et al. (2024b). In the emotional field, MuMA-ToM Shi et al. (2024) is used to evaluate the ability of agents to understand and reason about human interactions in a real home environment through video and text descriptions.
Training on generative agents. Li et al. (2024c) enhance the data to Supervised Fine-Tuning (SFT) generative agents with LLM-MASXu et al. (2023c) have created generative agents to overcome the intrinsic bias from LLMs by proposing a novel framework that powers generative agents with multi-agent reinforcement learning Xu et al. (2023c). For LLM-MAS, Yue et al. (2024) split complex trajectories in knowledge-intensive tasks into subtasks, proposing a co-training paradigm of the multi-agent framework, Long- and Short-Trajectory Learning, which ensures synergy while keeping the fine-grained performance of each agent. RLHF has been criticized for its high cost. Liu et al. (2023a) propose an alignment scheme based on a multi-agent system, effectively addressing instability and reward gaming concerns associated with reward-based RL optimization. Either way, LLM-MAS are essentially viewed as an environment in RL with different ways of getting rewards from the environment.
5.2 Resources of LLM-MAS for evaluations